# Classification and stability of penalized pinned elasticae

### v1

@@ 4. Energy comparison
We exhibit some elementary properties of $e$ and $h$ in the following lemmas, whose proof is postponed to Appendix B.
@@
[…]
@@
where $f$ is the function defined by (1.2). In particular, $h$ is decreasing on $(\frac{1}{\sqrt{2}},\hat{q}]$ and increasing on $[\hat{q},1)$.
Recall that we interpret the number $n\in\mathbf{N}$ as a mode. Therefore, as in [25], it will be naturally expected that, for instance, the energy $\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,n}]$ gets larger if $n$ is larger. However, this does not directly follow since the moduli $q=q_{i,n}$ depend on the choice of $n$, cf. Definition 2.7. In fact, combining the formula
@@
Hereafter we prove that $\Phi(\lambda)>0$ for any $\lambda\in(0,\ell^{-2}\hat{\lambda})$ and $\Phi(\ell^{-2}\hat{\lambda})=0$. Denote $q_{1}=q_{1,1}(\lambda):=q_{1}(\lambda\ell^{2})$ and $q_{2}=q_{2,1}(\lambda):=q_{2}(\lambda\ell^{2})$ for short. By Lemma 4.1, and the fact that $\lambda=\frac{1}{\ell^{2}}g(q_{1})=\frac{1}{\ell^{2}}g(q_{2})$, we deduce that
@@
Since $\mathrm{K}(q)$ and $(2\mathrm{E}(q)-\mathrm{K}(q))^{-1}$ are positive and strictly increasing with respect to $q\in(0,q_{*})$ (cf. Appendix A), we have $\Phi^{\prime}(\lambda)<0$ for $\lambda\in(0,\ell^{-2}\hat{\lambda})$. Moreover, $\Phi(\ell^{-2}\hat{\lambda})=0$ also follows since $\lambda=\ell^{-2}\hat{\lambda}$ implies that $q_{1}=q_{2}=\hat{q}$. Thus we obtain (4.13), and equality holds if and only if $\Phi(\lambda)=0$, i.e., $\lambda=\ell^{-2}\hat{\lambda}$.
@@
If $\lambda>0$ and $\ell>0$ satisfy $\lambda\ell^{2}<\hat{\lambda}$, then
$\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]<\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}].$
@@
We deduce from (4.6) that $\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]=2\sqrt{2}\sqrt{\lambda}h(q_{2,1})$ and $\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]=2\sqrt{2}\sqrt{\lambda}h(q_{3,1})$. Lemma 4.3 and the fact that $\hat{q}<q_{2,1}<q_{*}<q_{3,1}$ imply that $h(q_{2,1})<h(q_{3,1})$, so that $\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]<\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]$. ∎
@@
Since $q_{2}(\tfrac{\lambda\ell^{2}}{n^{2}})<q_{*}<q_{3}(\lambda\ell^{2})$, $\Psi_{\ell,n}^{\prime}(\lambda)$ takes a negative value for each $\lambda\in(0,n^{2}\ell^{-2}\hat{\lambda})$. The claim follows. ∎
@@
In view of Lemma 4.8, it suffices to show that $\Psi_{\ell,n}(n^{2}\ell^{-2}\hat{\lambda})>0$. By definition $q_{2}(\frac{\lambda\ell^{2}}{n^{2}})=\hat{q}$ holds when $\lambda=n^{2}\ell^{-2}\hat{\lambda}$. Throughout this proof write $\rho_{n}:=q_{3}(n^{2}\hat{\lambda})$ for short. Then, the problem is reduced to the positivity of $\Psi_{\ell,n}(n^{2}\ell^{-2}\hat{\lambda})=nh(\hat{q})-h(\rho_{n})$. Since $\hat{\lambda}=g(\hat{q})=n^{-2}g(\rho_{n})$ yields (after taking square roots) that
(4.17) $\sqrt{2\rho_{n}^{2}-1}\mathrm{K}(\rho_{n})=2\sqrt{2\rho_{n}^{2}-1}\mathrm{E}(\rho_{n})+n\sqrt{2\hat{q}^{2}-1}(2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q})),$
@@
$\frac{2\mathrm{E}(\rho_{n})-\mathrm{K}(\rho_{n})}{\sqrt{2\rho_{n}^{2}-1}}=-\frac{n\sqrt{2\hat{q}^{2}-1}}{2\rho_{n}^{2}-1}\big{(}2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q})\big{)}<-n\sqrt{2\hat{q}^{2}-1}\big{(}2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q})\big{)},$
@@
$\Psi_{\ell,n}(n^{2}\ell^{-2}\hat{\lambda})$ $=n\Big{(}2\sqrt{2\hat{q}^{2}-1}\mathrm{K}(\hat{q})+\frac{2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q})}{\sqrt{2\hat{q}^{2}-1}}\Big{)}-h(\rho_{n})$
$>3n\sqrt{2\hat{q}^{2}-1}\mathrm{K}(\hat{q})+n\frac{2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q})}{\sqrt{2\hat{q}^{2}-1}}-4\mathrm{E}(\hat{q})-2n\sqrt{2\hat{q}^{2}-1}\mathrm{E}(\hat{q})$
$=\frac{1}{\sqrt{2\hat{q}^{2}-1}}\Big{(}2n(3\hat{q}^{2}-2)\mathrm{K}(\hat{q})+4\big{(}n(1-\hat{q}^{2})-\sqrt{2\hat{q}^{2}-1}\big{)}\mathrm{E}(\hat{q})\Big{)}.$
Set $d_{n}:=2n(3\hat{q}^{2}-2)\mathrm{K}(\hat{q})+4(n(1-\hat{q}^{2})-\sqrt{2\hat{q}^{2}-1})\mathrm{E}(\hat{q})$ and hereafter we show that $d_{n}>0$. Since $\hat{q}$ is a solution of $f(q)=0$, $\hat{q}$ satisfies $(-4\hat{q}^{4}+5\hat{q}^{2}-1)\mathrm{K}(\hat{q})=(-8\hat{q}^{4}+8\hat{q}^{2}-1)\mathrm{E}(\hat{q})$. This yields
$d_{n}$ $=2n(3\hat{q}^{2}-2)\frac{-8\hat{q}^{4}+8\hat{q}^{2}-1}{-4\hat{q}^{4}+5\hat{q}^{2}-1}\mathrm{E}(\hat{q})+4\big{(}n(1-\hat{q}^{2})-\sqrt{2\hat{q}^{2}-1}\big{)}\mathrm{E}(\hat{q})$
$=2n\frac{-16\hat{q}^{6}+22\hat{q}^{4}-7\hat{q}^{2}}{-4\hat{q}^{4}+5\hat{q}^{2}-1}\mathrm{E}(\hat{q})-4\sqrt{2\hat{q}^{2}-1}\mathrm{E}(\hat{q}).$
Note that $-4\hat{q}^{4}+5\hat{q}^{2}-1=(4\hat{q}^{2}-1)(1-\hat{q}^{2})>0$ since $\hat{q}>{1}/{\sqrt{2}}$. By Lemma 2.5 and the fact that $[\frac{3}{5},\frac{2}{3}]\ni m\mapsto-16m^{3}+22m^{2}-7m$ is a positive function, we also find that $-16\hat{q}^{6}+22\hat{q}^{4}-7\hat{q}^{2}>0$ and hence
$d_{n}\geq d_{2}=\frac{1}{-4\hat{q}^{4}+5\hat{q}^{2}-1}\Big{(}4(-16\hat{q}^{6}+22\hat{q}^{4}-7\hat{q}^{2})-4\sqrt{2\hat{q}^{2}-1}(-4\hat{q}^{4}+5\hat{q}^{2}-1)\Big{)}\mathrm{E}(\hat{q}).$
@@
Notice that $m\mapsto-32m^{3}+36m^{2}-5m-3$ is strictly increasing in $[\tfrac{3}{5},\tfrac{2}{3}]$. Indeed, the derivative $m\mapsto-96m^{2}+72m-5$ has roots $m_{1,2}=\frac{3\pm\sqrt{\frac{17}{3}}}{8}$, implying that $m_{1}<0$ and $m_{2}>\frac{2}{3}$. Due to this monotonicity we infer $\psi(m)\geq-32(\tfrac{3}{5})^{3}+36(\tfrac{3}{5})^{2}-5\cdot\tfrac{3}{5}-3=\frac{6}{125}>0$. The positivity of $\Psi_{\ell,n}(n^{2}\ell^{-2}\hat{\lambda})$ follows and the proof is complete. ∎
@@
Let $\lambda>0$ and $\ell>0$ satisfy $\lambda\ell^{2}<\hat{\lambda}$. Then, there exists a unique $\lambda_{\dagger}\in(0,\ell^{-2}\hat{\lambda})$ such that
$\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]>\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]\ \ \text{if}\ \ \lambda<\lambda_{\dagger}\quad\text{and}\quad\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]<\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]\ \ \text{if}\ \ \lambda>\lambda_{\dagger}.$
In addition, $\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]=\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]$ if $\lambda=\lambda_{\dagger}$.
@@
Fix $\ell>0$ and define $\phi:(0,\ell^{-2}\hat{\lambda}]\to\mathbf{R}$ by
$\phi(\lambda):=\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]-\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}].$
Then, the fact that $\gamma_{\rm sarc}^{\hat{\lambda},\ell,1}=\gamma_{\rm larc}^{\hat{\lambda},\ell,1}$ and Lemma 4.6 yield that $\phi(\hat{\lambda})=\mathcal{E}_{\hat{\lambda}}[\gamma_{\rm larc}^{\hat{\lambda},\ell,1}]-\mathcal{E}_{\hat{\lambda}}[\gamma_{\rm loop}^{\hat{\lambda},\ell,1}]<0$. Moreover, noting that $q_{1}(\lambda\ell^{2})\to 1/\sqrt{2}$ and $q_{3}(\lambda\ell^{2})\to q_{*}$ as $\lambda\to 0$, we deduce from the energy formulae (4.1) that
$\lim_{\lambda\to 0}\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]$ $=\frac{8}{\ell}\big{(}2\mathrm{E}(\tfrac{1}{\sqrt{2}})-\mathrm{K}(\tfrac{1}{\sqrt{2}})\big{)}^{2}>0,$
$\lim_{\lambda\to 0}\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]$ $=0,$
from which it follows that $\phi(0+)>0$. Thus, for the desired conclusion it suffices to check that $\phi$ is strictly decreasing on $(0,\ell^{-2}\hat{\lambda}]$. As in (4.14), we compute
$\phi^{\prime}(\lambda)=$ $\,\frac{dq}{d\lambda}\Big{|}_{q=q_{1}}\frac{d}{dq}\Big{(}8|2\mathrm{E}(q)-\mathrm{K}(q)|\big{(}(4q^{2}-3)\mathrm{K}(q)+2\mathrm{E}(q)\big{)}\Big{)}\Big{|}_{q=q_{1}}$
@@
Since $\lambda\ell^{2}=g(q_{1}(\lambda\ell^{2}))=g(q_{3}(\lambda\ell^{2}))$ yield $2\mathrm{E}(q_{1})-\mathrm{K}(q_{1})=\frac{1}{2\sqrt{2}}\sqrt{\lambda}\ell(2q_{1}^{2}-1)^{-\frac{1}{2}}$ and $\mathrm{K}(q_{3})-2\mathrm{E}(q_{3})=\frac{1}{2\sqrt{2}}\sqrt{\lambda}\ell(2q_{3}^{2}-1)^{-\frac{1}{2}}$, we obtain
$\phi^{\prime}(\lambda)=\frac{2\sqrt{2}\ell}{\sqrt{\lambda}}\bigg{(}\mathrm{K}(q_{1})\sqrt{2q_{1}^{2}-1}-\mathrm{K}(q_{3})\sqrt{2q_{3}^{2}-1}\bigg{)}.$
Then, in view of the fact that the map $q\mapsto\mathrm{K}(q)\sqrt{2q^{2}-1}$ is strictly increasing, we find that the right-hand side in the above equation takes a negative value for all $\lambda\in(0,\hat{\lambda}\ell^{-2})$. The proof is complete. ∎
@@
  • $\gamma_{\rm sarc}^{\lambda,\ell,1},\gamma_{\rm loop}^{\lambda,\ell,1}$ have a larger energy than $\gamma_{\rm larc}^{\lambda,\ell,1}$, but their order depends on $\lambda$, cf. Lemma 4.10. As seen in Figure 4 the order changes once at $\lambda_{\dagger}\simeq 0.32241$.
@@
Figure 4. Plots of $\lambda\mapsto\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,1,1}]$ (blue), $\lambda\mapsto\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,1,1}]$ (yellow) and $\lambda\mapsto\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,1,1}]$ (green) in the for $\lambda\in(0.02,0.7)$

### v2

@@ 4. Energy comparison
We exhibit some elementary properties of $e$ and $h$ in the following lemmas, whose proofs are postponed to Appendix B.
@@
(4.8) $\text{ strictly decreasing in }(\tfrac{1}{\sqrt{2}},\hat{q})\cup(q_{*},1),\quad\text{ strictly increasing in }(\hat{q},q_{*}).$
@@
where $f$ is the function defined by (1.2). In particular, $h$ is decreasing on $(\frac{1}{\sqrt{2}},\hat{q}$] and increasing on $\hat{q},1)$.
Recall that we interpret the number $n\in\mathbf{N}$ as a mode. Therefore, as in [Ydcds], it will be naturally expected that, for instance, the energy $\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,n}]$ gets larger if $n$ is larger. However, this does not directly follow since the moduli $q=q_{i,n}$ depend on the choice of $n$, cf. Definition 2.7. In fact, combining the formula
@@
[…]
@@
[…]
@@
If $\lambda>0$ and $\ell>0$ satisfy $\lambda\ell^{2}\leq\hat{\lambda}$, then
$\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]<\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}],$
where the left-hand side is interpreted as $\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]=\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]$ if $\lambda\ell^{2}=\hat{\lambda}$.
@@
We deduce from (4.6) that $\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]=2\sqrt{2}\sqrt{\lambda}h(q_{2,1})$ and $\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]=2\sqrt{2}\sqrt{\lambda}h(q_{3,1})$. Lemma 4.3 and the fact that $\hat{q}\leq q_{2,1}<q_{*}<q_{3,1}$ imply that $h(q_{2,1})<h(q_{3,1})$, so that $\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]<\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]$. ∎
@@
[…]
@@
[…]
(4.17) $),\displaystyle\sqrt{2\rho_{n}^{2}-1}\mathrm{K}(\rho_{n})=2\sqrt{2\rho_{n}^{2}-1}\mathrm{E}(\rho_{n})+n\sqrt{2\hat{q}^{2}-1}(2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q}$)),
@@
$),\frac{2\mathrm{E}(\rho_{n})-\mathrm{K}(\rho_{n})}{\sqrt{2\rho_{n}^{2}-1}}=-\frac{n\sqrt{2\hat{q}^{2}-1}}{2\rho_{n}^{2}-1}\big(2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q})\big)<-n\sqrt{2\hat{q}^{2}-1}\big(2\mathrm{E}(\hat{q})-\mathrm{K}(\hat{q}$$\big)$,
@@
[…]
$).\displaystyle=\frac{1}{\sqrt{2\hat{q}^{2}-1}}\Big(2n(3\hat{q}^{2}-2)\mathrm{K}(\hat{q})+4\big(n(1-\hat{q}^{2})-\sqrt{2\hat{q}^{2}-1}\big)\mathrm{E}(\hat{q}$$\Big)$.
[…]
Note that $4\hat{q}^{4}+5\hat{q}^{2}-1=(4\hat{q}^{2}-1)(1-\hat{q}^{2})>0$ since $\hat{q}>{1}/{\sqrt{2}}$. By Lemma 2.5 and the fact that $\frac{3}{5},\frac{2}{3}]\ni m\mapsto-16m^{3}+22m^{2}-7m$ is a positive function, we also find that $16\hat{q}^{6}+22\hat{q}^{4}-7\hat{q}^{2}>0$ and hence
[…]
@@
Notice that $m\mapsto-32m^{3}+36m^{2}-5m-3$ is strictly increasing in $\tfrac{3}{5},\tfrac{2}{3}]$. Indeed, the derivative $m\mapsto-96m^{2}+72m-5$ has roots $m_{1,2}=\frac{3\pm\sqrt{\frac{17}{3}}}{8}$, implying that $m_{1}<0$ and $m_{2}>\frac{2}{3}$. Due to this monotonicity we infer $\psi(m)\geq-32(\tfrac{3}{5})^{3}+36(\tfrac{3}{5})^{2}-5\cdot\tfrac{3}{5}-3=\frac{6}{125}>0$. The positivity $\Psi_{\ell,n}(n^{2}\ell^{-2}\hat{\lambda}) fo$llows and the proof is complete. ∎
@@
[…]
$\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]>\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]\ \ \text{if}\ \ \lambda\ell^{2}<\lambda_{\dagger}\quad\text{and}\quad\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]<\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]\ \ \text{if}\ \ \lambda\ell^{2}>\lambda_{\dagger}.$
In addition, $\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]=\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]$ if $\lambda\ell^{2}=\lambda_{\dagger}$.
@@
For $\ell>0$ we define $\phi_{\ell}:(0,\hat{\lambda}]\to\mathbf{R}$ by
$\phi_{\ell}(\lambda):=\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]-\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}].$
From (4.1) (and the fact that $q_{i,1}=q_{i}(\lambda\ell^{2})$ for $i=1,3$) one infers that
(4.18) $\phi_{\ell}(\lambda)=\frac{1}{\ell}\phi_{1}(\lambda\ell^{2}).$
Since the energy order depends only on the sign of $\phi_{\ell}$ (and thus of $\phi_{1}(\cdot\ell^{2})$) the above formula allows us to infer the claim from the study of the special case of $\ell=1$. The fact that $\gamma_{\rm sarc}^{\hat{\lambda},1,1}=\gamma_{\rm larc}^{\hat{\lambda},1,1}$ and Lemma 4.6 yield that $\phi_{1}(\hat{\lambda})=\mathcal{E}_{\hat{\lambda}}[\gamma_{\rm larc}^{\hat{\lambda},1,1}]-\mathcal{E}_{\hat{\lambda}}[\gamma_{\rm loop}^{\hat{\lambda},1,1}]<0$. Moreover, noting that $q_{1}(\lambda)\to 1/\sqrt{2}$ and $q_{3}(\lambda)\to q_{*}$ as $\lambda\to 0$, we deduce from the energy formulae (4.1) that
$\lim_{\lambda\to 0}\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,1,1}]$ $=8\big(2\mathrm{E}(\tfrac{1}{\sqrt{2}})-\mathrm{K}(\tfrac{1}{\sqrt{2}})\big)^{2}>0,$
$\lim_{\lambda\to 0}\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,1,1}]$ $=0,$
from which it follows that $\phi_{1}(0+)>0$. Next we show that $\phi_{1}$ is strictly decreasing on $(0,\hat{\lambda}$]. As in (4.14), we compute
$\phi_{1}^{\prime}(\lambda)=$ $\,\frac{dq}{d\lambda}\Big|_{q=q_{1}}\frac{d}{dq}\Big(8|2\mathrm{E}(q)-\mathrm{K}(q)|\big((4q^{2}-3)\mathrm{K}(q)+2\mathrm{E}(q)\big)\Big)\Big|_{q=q_{1}}$
@@
Since $\lambda=g(q_{1}(\lambda))=g(q_{3}(\lambda))$ yield $2\mathrm{E}(q_{1})-\mathrm{K}(q_{1})=\frac{1}{2\sqrt{2}}\sqrt{\lambda}(2q_{1}^{2}-1)^{-\frac{1}{2}}$ and $\mathrm{K}(q_{3})-2\mathrm{E}(q_{3})=\frac{1}{2\sqrt{2}}\sqrt{\lambda}(2q_{3}^{2}-1)^{-\frac{1}{2}}$, we obtain
$\phi_{1}^{\prime}(\lambda)=\frac{2\sqrt{2}}{\sqrt{\lambda}}\bigg(\mathrm{K}(q_{1})\sqrt{2q_{1}^{2}-1}-\mathrm{K}(q_{3})\sqrt{2q_{3}^{2}-1}\bigg).$
[…]
$\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,\ell,1}]-\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]=\phi_{\ell}(\lambda)=\frac{1}{\ell}\phi_{1}(\lambda\ell^{2})>\frac{1}{\ell}\phi_{1}(\lambda_{\dagger})=0.$
Similarly the remaining assertions follow. The proof is complete. ∎
@@
  • $\gamma_{\rm sarc}^{\lambda,\ell,1},\gamma_{\rm loop}^{\lambda,\ell,1}$ have a larger energy than $\gamma_{\rm larc}^{\lambda,\ell,1}$, but their order depends on $\lambda$, cf. Lemma 4.10. More precisely we have shown that the order changes once at $\ell^{-2}\lambda_{\dagger}$. From Figure 4, which shows the special case of $\ell=1$, one can read off that $\lambda_{\dagger}\simeq 0.32241$.
@@
Figure 4. Plots of $\lambda\mapsto\mathcal{E}_{\lambda}[\gamma_{\rm sarc}^{\lambda,1,1}]$ (blue), $\lambda\mapsto\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,1,1}]$ (yellow) and $\lambda\mapsto\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,1,1}]$ (green) for $\lambda\in(0.02,0.7)$